You are told that Lands End, a catalog retailer, earned an excess return (Jensens alpha), in annualized terms, of 32% over
Answer #1

If we put the given information in regression equation.

Stock returns = Rs = 32%

Market Return = Rm = 4.8%

The beta (β) is = \beta = 1.5

Rs = \alpha + \beta *R_{m}

32% = \alpha + 1.5* 4.8

32% = \alpha + 7.2 %

\alpha = 32%/7.2%%/ 7.2%

4.44% is the intercept.

Hope this helps! Feel free to share your feedback. Thanks and have a good day.